The footings A and B carry the weight of the silo. Due to this load, the
probability of excessive settlement in each footing is 0.17. If footing A is settled, the probability
of the settlement of the footing B is 0,8. When footing B is settled, the probability of the
settlement of the footing A is 0,87. What is the probability of the differential settlement.

Answer by ikleyn(52908)

.

    P = P(A is settled, B is not) + P(A is not settled, B is) = 0.17*(1-0.8) + 0.17*(1-0.87) = 0.0561.   ANSWER

Solved.